OpenStudy (anonymous):

Aileen deposited money into an account in which interest is compounded semiannually at a rate of 2.1%. How much did she deposit if the total amount in her account after 3 years was $3952.08, and she made no other deposits or withdrawals? (Points : 3) $3489 $3562 $3712 $3751

OpenStudy (anonymous):

please help @amistre64

OpenStudy (amistre64):

what have you tried?

OpenStudy (anonymous):

thats the thing i dont know what to try ive nevr done something like this because i have been behind in school and wasnt there to see the lesson, i dont know what to start with :(

OpenStudy (amistre64):

start with the basics, what does interest do?

OpenStudy (anonymous):

adds more?

OpenStudy (amistre64):

yep, we we add 2.1% more to the account each period.

OpenStudy (amistre64):

how do you think we would calculate that increase?

OpenStudy (anonymous):

hmmm do we add it to the big number>

OpenStudy (amistre64):

if i have $100, and interest is 2%, what is the amount of interest that was earned in one interest period? 100(.02) = 2, the interest earned is $2 right?

OpenStudy (anonymous):

the 3952.08 number

OpenStudy (amistre64):

lets work on the process, since you missed a lesson on it

OpenStudy (anonymous):

yes its 2 % of your original number

OpenStudy (amistre64):

so, when we compound, we add the interest back to our account balance. 100 + 100(.02) 100(1 + .02) 100(1.02) = 102 does this make sense?

OpenStudy (anonymous):

yes i think so, your plugging in numbers to get the 102

OpenStudy (amistre64):

yeah now, the next period we have interest that is calculated on 102 because thats our NEW balance in the bank 102 + 102(.02) 102(1.02) but we really want to view this from our original balance, recall that 102 = 100(1.02) soo 100(1.02)(1.02 ) = 104.04 or written more traditionally 100(1.02)^2 = 104.04

OpenStudy (amistre64):

if we compound it again for another period, what do you think our calculation will be in terms of the original balance? do you see a pattern?

OpenStudy (anonymous):

um i do but i dont think its relevant to the problem

OpenStudy (amistre64):

why would i speak of things that arent relevant to the problem?

OpenStudy (amistre64):

what is the pattern that you see?

OpenStudy (anonymous):

well i see that there is an increase of 2. firt was 102.2 then 104.4, its like 2 4 ..6 ...8 and so on

OpenStudy (anonymous):

no i wasnt saying u would saying it was irrelevant i was saying my answer was

OpenStudy (amistre64):

oh, lol ... at least your trying, thats importnat, ill guide you back on track with any luck

OpenStudy (amistre64):

the pattern is that we have a starting balance of 100, and an interest that compounds at (1.02) and the exponent on the interest is equal to the period we are trying to find. say the 5th period, what is our balance? 100 (1.02)^5 = 110.408... or 110.41 theres only one other thing we need to do to solve your problem, maybe 2 things

OpenStudy (amistre64):

lets see first if you understand this concept what is our 7th period balance? what would we setup?

OpenStudy (anonymous):

gaaaawwwww im so confused lol sorry im trying to understand but im lost

OpenStudy (amistre64):

the pattern is in changing the exponent ... the 3rd period, what is our balance? ^^^^^^^ 100 (1.02)^(3) = 106.120.. or 106.12 ^^ ------------------------------ the 5th period, what is our balance? ^^^^^^^ 100 (1.02)^(5) = 110.408... or 110.41 ^^ ------------------------------ the 27th period, what is our balance? ^^^^^^^ 100 (1.02)^(27) = 170.688... or 170.69 ^^ ----------------------------- all that changes is an exponent

OpenStudy (amistre64):

now try to tell me, show me your work, on how we find the 7th period balance.

OpenStudy (anonymous):

okay 100(1.02)^(7)

OpenStudy (amistre64):

perfect ..... thats all there is to the calculation let put this into a formula, so that all we have to do is plug in values, \[A=P(1+r)^n\] P is, Principle amount (starting balance) A is, the balance Amount we have at a given period r is the interest rate for that period does this make sense?

OpenStudy (anonymous):

yes i think lol

OpenStudy (amistre64):

its exposure at least, we will use this formula to solve your problem, but we need to define r and n to do it. your problem says it compounds semiannually - which simply means it occurs 2 times a year. if something happens 2 times a year, for 3 years. how many times has it occured?

OpenStudy (anonymous):

6

OpenStudy (amistre64):

good, then n=6 now we need the proper rate adjustment. we are given a yearly interest rate of 2.1%, we need to DIVIDE this over the year in order to determine the 'period' interest rate. 2 times a year, wouldnt we divide 2.1 in half? that way in 1 year we have used the full 2.1% value. tell me what im saying for the interest rate

OpenStudy (anonymous):

well would you 2.1 to 2.1 cause its 2 times a year?

OpenStudy (amistre64):

lets say we have $2.10 to spread out over 2 people. how much does each person get?

OpenStudy (anonymous):

1.05

OpenStudy (amistre64):

perfect you agree that its fair to give each person $1.05. (2.10)/2 = 1.05 same concept applies here, we want to spread out 2.10% over 2 periods in the year. each period therefore gets 1.05% r = 1.05%, or written in decimal form: r = .0105

OpenStudy (amistre64):

we are given an ending balance of: A = 3952.08 n = 6, and r = .0105 \[A=P(1+r)^n\] \[3952.08=P(1+\frac{.0210}{2})^{2*3}\] \[3952.08=P(1+.0105)^{6}\] \[3952.08=P(1.0105)^{6}\] how do we solve for P?

OpenStudy (anonymous):

do the (1.0105)^6

OpenStudy (amistre64):

we could, but eventually what do we need to do? divide right? n = jm, then to solve for j j = n/m

OpenStudy (amistre64):

the numbers are going to be what they are, but the process remains true regardless

OpenStudy (amistre64):

\[P=\frac{A}{(1+r)^n}\]

OpenStudy (amistre64):

P = 3952.08/(1.0105)^6 all that background, just for that lol

OpenStudy (anonymous):

ok okay lol i see hahah sorry

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

$3712 is the answer

OpenStudy (amistre64):

yes, yes it is